OpenStudy (anonymous):

lim x goes to infiniti, (x+1)^((1/(ln(x+1))

OpenStudy (anonymous):

according to wolframalpha, the answer is e, but i cant figure out how

OpenStudy (anonymous):

use l'hopital 's rule and it will eventually be e

OpenStudy (anonymous):

\[\lim_{x\rightarrow\infty}(1+x)^{\frac{1}{\ln(x+1)}}\]

OpenStudy (anonymous):

step one is as usual take the log

OpenStudy (anonymous):

you get \[\frac{1}{\ln(x+1)}\times \ln(x+1)\]

OpenStudy (anonymous):

first take down the exponent, and put it like 1/(ln(x+1)) * (ln(x+1))

OpenStudy (anonymous):

that limit is pretty obviously one, since it is identically one

OpenStudy (anonymous):

but I got e on my CAS software

OpenStudy (anonymous):

so your answer is \[e^1=e\]

OpenStudy (anonymous):

ooooh that is right :-D

OpenStudy (anonymous):

wait the answer is e, not 1 step one was to take the log step two is compute the limits step three is exponentiate since step one was take the log

OpenStudy (anonymous):

huh? im confused with step one? do you mind breaking it down into baby steps?

OpenStudy (anonymous):

ok... foo! lol... I'm determined .. ha ha... \[y = \ (x+1)^{1/\ln (x+1)}\] \[\ln y = \ln (x+1)^{1/\ln (x+1)}\] \[\ln y = (1/\ln (x+1))(\ln (x+1))\] \[e ^{\ln y} = e ^{\ln (x+1)/\ln (x+1)}\] y = e^1 dy/dx = e^1 = e

OpenStudy (anonymous):

where did you get the y from?